On a snowy day a teen goes tobogganing. The teen pushes the tobbogan with an initial velocity of 1.0 m/s and it begins to slide down a bumpy hill as shown. If the height of the hill is 15 m what is the velocity of the toboggan at the bottom of said hill? Ignore friction.
Answer #1

energy conservation

K.E. (INITIAL ) + P.E. (INITIAL ) = K.E. (FINAL)

(1/2) M U² + MgH = (1/2) MV²

(1/2 ) U² + gH = (1/2) V²

(0.5) (1)² + 9.8 ( 15) = (0.5) V²

0.5 + 147 = 0.5 V²

V² = 147.5 / 0.5 = 295

V = 17.175 m/s